# Data-Efficient Learning with Neural Programs

### v1

@@ Abstract
Many computational tasks can be naturally expressed as a composition of a DNN followed by a program written in a traditional programming language or an API call to an LLM. We call such composites "neural programs" and focus on the problem of learning the DNN parameters when the training data consist of end-to-end input-output labels for the composite. When the program is written in a differentiable logic programming language, techniques from neurosymbolic learning are applicable, but in general, the learning for neural programs requires estimating the gradients of black-box components. We present an algorithm for learning neural programs, called ISED, that only relies on input-output samples of black-box components. For evaluation, we introduce new benchmarks that involve calls to modern LLMs such as GPT-4 and also consider benchmarks from the neurosymolic learning literature. Our evaluation shows that for the latter benchmarks, ISED has comparable performance to state-of-the-art neurosymbolic frameworks. For the former, we use adaptations of prior work on gradient approximations of black-box components as a baseline, and show that ISED achieves comparable accuracy but in a more data- and sample-efficient manner.  [^1] Code is available at https://github.com/alaiasolkobreslin/ISED/tree/v1.0.0
@@ 1 Introduction
Many computational tasks cannot be solved by neural perception alone but can be naturally expressed as a composition of a neural model $M_{\theta}$ followed by a program $P$ written in a traditional programming language or an API call to an LLM. We call such composites "neural programs" and study the problem of learning neural programs in an end-to-end manner with a focus on data and sample efficiency. One problem that is naturally expressed as a neural program is scene recognition [29], where $M_{\theta}$ classifies objects in an image and $P$ prompts GPT-4 to identify the room type given these objects (Fig. 1).
@@
Learning neural programs when $P$ is not expressed as a logic program is a difficult problem because gradients across black-box programs cannot be computed explicitly. One possible solution is to use REINFORCE [26] to sample symbols from distributions predicted by $M_{\theta}$ and compute the expected reward using the output label. However, REINFORCE is not sample-efficient as it produces a weak learning signal, especially when applied to programs with a large number of inputs. There are other REINFORCE-based methods that can be applied to the neural program learning setting, namely IndeCateR [21] and Neural Attention for Symbolic Reasoning (NASR) [5]. However, IndeCateR struggles with sample efficiency despite providing lower variance than REINFORCE, and NASR performs poorly when intermediate labels are unavailable for pretraining. Another possible solution is Approximate Neurosymbolic Inference (A-NeSI) [24], which trains a neural model to estimate the gradient of $P$, but learning the surrogate neural network becomes more difficult as the complexity of $P$ increases. Moreover, the additional neural models in the learning framework in A-NeSI results in data inefficiency.
@@
Our evaluation considers 16 neural program benchmark tasks. Our results show that ISED outperforms purely neural networks and CLIP [19] on new neural program tasks involving GPT-4 calls. Additionally, ISED outperforms neurosymbolic methods on 9 of the 14 benchmarks tasks that can be encoded in logic programming languages. ISED is also the top performer on 8 out of the 16 benchmark tasks when compared to REINFORCE-based and black-box gradient estimation baselines. Furthermore, we show that ISED is more data- and sample-efficient than baseline methods.
@@ 2 Neural Programs
With ISED, the program can be written in just a few lines of Python. It takes in a list of characters representing symbols, and simply invokes the Python eval function on the joined expression string. Under ISED, the hwf evaluation function can be used just like any other PyTorch [17] module, and will internally perform sampling and probability estimation to facilitate learning.
@@ 3.1 ISED Overview
Assuming $P$ is a black-box, we can collect symbol-output samples $(u,v)$ from $P$. Such collection of samples can be viewed as a summary logic program consisting of rules of the form if $r=u$ then $y=v$. For instance, in the task of adding two digits $r_{1}$ and $r_{2}$, one rule of the logic program would be $r_{1}=1\land r_{2}=2\rightarrow y=3$. Techniques from neurosymbolic literature via exact or approximate weighted model counting (WMC) [10] can then be used for computing the gradient across such a summary of $P$. However, having the complete summary of all combinations of symbols is not feasible for a black-box $P$. ISED samples symbols from the probability distribution predicted by the neural network $M_{\theta}$, evaluates $P$ on each sample, and takes the gradient across this partial summary of $P$. This is a good approximation of the complete summary since it is likely to contain symbols with high probability which contribute the most in exact computation.
This approach differs from REINFORCE in how it differentiates through this summary of $P$. REINFORCE rewards sampled symbols that resulted in the correct output through optimizing the log probability of the samples, weighted by reward values. This weighted-sum style estimation provides a weaker learning signal compared to WMC used by ISED, making learning harder for REINFORCE as the number of inputs to $P$ increases. See Appendix A for further details.
@@ 3.2 Preliminaries and Programming Interface
Using this, we may further define data-types such as $\textsc{Integer}_{j}^{k}=\textsc{Discrete}(\{j,\dots,k\})$, $\textsc{Digit}=\textsc{Integer}_{0}^{9}$, and $\textsc{Bool}=\textsc{Discrete}(\{\text{true},\text{false}\})$.
We also define a black-box program $P$ as a function $(\tau_{1},\dots,\tau_{m})\rightarrow\tau_{o}$, where $\tau_{1},\dots,\tau_{m}$ are the input types and $\tau_{o}$ is the output type. For example, the structural input mapping for the hand-written formula task is $\textsc{List}_{7}(\textsc{Discrete}(\{0,\dots,9,+,-,\times,\div\}))$, and the structural output mapping is Float. The mapping suggests that the program takes a list of length up to 7 as input, where each element is a digit or an arithmetic operator, and returns a floating point number.
@@
In order to represent the ground truth output as a distribution to be used in the loss function, there needs to be a mechanism for transforming SET($\tau$) mappings into DIST($\tau$) mappings. For this purpose, we define a vectorize function $\delta_{\tau}:(\text{SET}(\tau),2^{\tau})\rightarrow\text{DIST}(\tau)$ for the different output mappings $\tau$ in Table 2. When considering a datapoint $(x,y)$ during training, ISED samples many symbols and obtains a list of outputs $\hat{y}$. The vectorizer then takes the ground truth $y$ and the outputs $\hat{y}$ as input and returns the equivalent distribution interpretation of $y$. While $\hat{y}$ is not used by $\delta_{\tau}$ in most cases, we include it as an argument so that Float output mappings can be discretized, which is necessary for vectorization. For example, if the inputs to the vectorizer for the hand-written formula task are $y=2.0$ and $\hat{y}=[1.0,3.5,2.0,8.0]$, then it would return $[0,0,1,0]$.
@@ 3.3 Algorithm
We now formally present the ISED algorithm. For a given task, there is a black-box program $P$, taking $m$ inputs, that operates on structured data. Let $\tau_{1},...,\tau_{m}$ be the mappings for these inputs and $\tau_{o}$ the mapping for the program’s output. We write $P$ as a function from its input mappings to its output mapping: $P:(\tau_{1},...,\tau_{m})\rightarrow\tau_{o}$. For each unstructured input $i$ to the program, there is a neural model $M^{i}_{\theta_{i}}:x_{i}\rightarrow\text{DIST}(\tau_{i})$. $S$ is a sampling strategy (e.g., categorical sampling) that samples symbols using the outputs of a neural model, and $k$ is the number of samples to take for each training example. There is also a loss function $\mathcal{L}$ whose first and second arguments are the prediction and target values respectively. We present the algorithm in pseudocode in Algorithm 1 and describe its steps with the hand-written formula task:
@@
13:         $\hat{w}\leftarrow\text{normalize}([\omega(y_{k},\hat{y},\hat{r},\hat{p})\text{for }y_{k}\in\tau_{o}\text{ (or }y_{k}\in\hat{y})])$triangleright Estimate
@@
Estimate. In order to compute the prediction value to use in the loss function, ISED must consider each output $y_{k}$ in the output mapping and accumulate the aggregated probabilities for all sampled symbols that resulted in output $y_{k}$. We specify $\otimes$ as the min function, and $\oplus$ as the max function in this example. Note that ISED requires that $\otimes$ and $\oplus$ represent either min and max or mult and add respectively. We refer to these two options as the min-max and add-mult semirings. We define an accumulate function $\omega$ that takes as input an element of the output mapping $y_{k}$, sampled outputs $\hat{y}$, sampled symbols $\hat{r}$, and predicted input distributions $\hat{p}$. The accumulator performs the $\oplus$ operation on aggregated probabilities for elements of $\hat{y}$ that are equal to $y_{k}$ and is defined as follows:
$\omega(y_{k},\hat{y},\hat{r},\hat{p})=\oplus_{j=1}^{k}\mathbf{1}_{\hat{y}_{j}=y_{k}}\sigma_{\tau_{o}}(\hat{r}_{j},\hat{p}_{j})$
ISED then sets $\tilde{w}=[\omega(y_{k},\hat{y},\hat{r},\hat{p})\text{ for }y_{k}\in\tau_{o}]$ in the case where $\tau_{o}$ is not Float. When $\tau_{o}$ is Float, as for hand-written formula, it only considers $y_{k}\in\hat{y}$. Next, it performs $L_{2}$ normalization over each element in $\tilde{w}$ and sets $\hat{w}$ to this result. To initialize the ground truth vector, it sets $w=\delta(y,\hat{y})$. ISED then initializes $l=\mathcal{L}(\hat{w},w)$ and computes $\frac{\partial l}{\partial\theta_{i}}$ for each input $i$. These steps are shown on lines 13-16.
@@ 4.1 Benchmark Tasks: NeuroGPT, NeuroPython, and Neurosymbolic
We also consider several tasks from the neurosymbolic literature, including Sudoku solving and hand-written formula (HWF) evaluation. While the solutions to many of these tasks are usually presented as a logic program in neurosymbolic learning frameworks, neural program solutions can take the form of Python programs. We call such models neuroPython programs.
@@
Visual Sudoku. The goal of this task is to solve an incomplete 9x9 Sudoku, where the problem board is given as MNSIT digits. We follow the experimental setting used in NASR, including their pre-trained MNIST digit recognition models [5]. We use the SatNet dataset consisting of 9K training samples and 500 test samples [25].
@@ 4.2 Evaluation Setup and Baselines
We pick as baselines neurosymbolic methods DeepProbLog (DPL) [14] and Scallop [13], A-NeSI [24] which performs neural approximation of the gradients, and sampling-based gradient approximation methods REINFORCE [26], IndeCateR [21], and NASR [5]. IndeCateR achieves provably lower variance than REINFORCE by using a specialized sampling method (Appendix A), and NASR is a variant specialized for efficient finetuning by using a single sample and a custom reward function. We also use purely neural baselines and CLIP [19] for GPT leaf and scene. CLIP is a multimodal model that supports zero-data image classification by simply providing names of the categories.
@@ 4.3 RQ1: Performance and Accuracy
Table 3: Performance on selected benchmarks. "TO" means time-out, and "N/A" means the task could not be programmed in the framework. Methods are divided (from top to bottom) by neurosymbolic, black-box gradient estimation, and REINFORCE-based.
@@
Method sum4 HWF DT leaf GPT leaf scene sudoku
DPL TO TO $39.70$ N/A N/A TO
Scallop $80.10$ $96.65$ $81.13$ N/A N/A TO
A-NeSI $78.10$ $3.13$ $78.82$ $72.40$ $61.46$ $26.36$
REINFORCE $13.84$ $18.59$ $23.60$ $34.02$ $47.07$ $79.08$
IndeCateR $13.02$ $15.14$ $40.38$ $52.67$ $12.28$ $66.50$
NASR $4.86$ $1.85$ $16.41$ $17.32$ $2.02$ $82.78$
ISED (ours) $94.10$ $97.34$ $82.32$ $79.95$ $68.59$ $80.32$
To answer RQ1, we evaluate ISED’s accuracy against those of the baselines. ISED matches, and in many cases surpasses, the accuracy of neurosymbolic and gradient estimation baselines. We highlight the results for sum4 from MNIST-R and other benchmarks in Table 3. Tables 7-10 in Appendix C contain results for the remaining MNIST-R tasks, including standard deviations for all tasks. ISED is the top performer on 8 out of the 16 total tasks. On the GPT leaf and scene tasks, ISED outperforms the purely neural baseline by $3.82\%$ and $31.42\%$ respectively, and zero-shot CLIP by $59.80\%$ and $17.50\%$. For many tasks, A-NeSI is the non-neurosymbolic method that comes closest to ISED, sometimes outperforming our method. However, A-NeSI achieves significantly lower performance than ISED on tasks involving complex programs, namely HWF and sudoku. This is likely due to the difficulty of training a neural model to estimate the output of $P$ and its gradient when $P$ is complex. ISED also outperforms REINFORCE on all but 2 tasks due to the REINFORCE learning signal being weaker for tasks where $P$ involves multiple inputs. NASR outperforms ISED only on sudoku by $2.46\%$ due to NASR being well-suited for fine-tuning as it restricts its algorithm to use one sample. ISED also outperforms IndeCateR on all but 4 tasks due to IndeCateR being less sample efficient than ISED.

### v2

@@ Abstract
Many computational tasks can be naturally expressed as a composition of a DNN followed by a program written in a traditional programming language or an API call to an LLM. We call such composites “neural programs” and focus on the problem of learning the DNN parameters when the training data consist of end-to-end input-output labels for the composite. When the program is written in a differentiable logic programming language, techniques from neurosymbolic learning are applicable, but in general, the learning for neural programs requires estimating the gradients of black-box components. We present an algorithm for learning neural programs, called ISED, that only relies on input-output samples of black-box components. For evaluation, we introduce new benchmarks that involve calls to modern LLMs such as GPT-4 and also consider benchmarks from the neurosymbolic learning literature. Our evaluation shows that for the latter benchmarks, ISED has comparable performance to state-of-the-art neurosymbolic frameworks. For the former, we use adaptations of prior work on gradient approximations of black-box components as a baseline, and show that ISED achieves comparable accuracy but in a more data- and sample-efficient manner.  [^1] Code is available at https://github.com/alaiasolkobreslin/ISED
@@ 1 Introduction
Many computational tasks cannot be solved by neural perception alone but can be naturally expressed as a composition of a neural model $M_{\theta}$ followed by a program $P$ written in a traditional programming language or an API call to a large language model (LLM). We call such composites “neural programs” and study the problem of learning neural programs in an end-to-end manner with a focus on data and sample efficiency. One problem that is naturally expressed as a neural program is scene recognition [29], where $M_{\theta}$ classifies objects in an image and $P$ prompts GPT-4 to identify the room type given these objects (Fig. 1).
@@
Learning neural programs when $P$ is not expressed as a logic program is a difficult problem because gradients across black-box programs cannot be computed explicitly. One possible solution is to use REINFORCE [26] to sample symbols from distributions predicted by $M_{\theta}$ and compute the expected reward using the output label. However, REINFORCE is not sample-efficient as it produces a weak learning signal, especially when applied to programs with a large number of inputs. There are other REINFORCE-based methods that can be applied to the neural program learning setting, namely IndeCateR [21] and Neural Attention for Symbolic Reasoning (NASR) [5]. However, IndeCateR struggles with sample efficiency despite providing lower variance than REINFORCE, and NASR performs poorly when intermediate labels are unavailable for pretraining. Another possible solution is Approximate Neurosymbolic Inference (A-NeSI) [24], which trains a neural network to estimate the gradient of $P$, but learning the surrogate neural network becomes more difficult as the complexity of $P$ increases. Moreover, the additional neural models in the learning framework in A-NeSI results in data inefficiency.
@@
Our evaluation considers 16 neural program benchmark tasks. Our results show that ISED outperforms purely neural networks and CLIP [19] on neural program tasks involving GPT-4 calls. Additionally, ISED outperforms neurosymbolic methods on 9 of the 14 benchmarks tasks that can be encoded in logic programming languages. ISED is also the top performer on 8 out of the 16 benchmark tasks when compared to REINFORCE-based and black-box gradient estimation baselines. Furthermore, we show that ISED is more data- and sample-efficient than baseline methods.
@@ 2 Neural Programs
With ISED, the program can be written in just a few lines of Python. It takes in a list of characters representing symbols, and simply invokes the Python eval function on the joined expression string. The hwf evaluation function can be used just like any other PyTorch [17] module since ISED internally performs sampling and probability estimation to estimate the gradient.
@@ 3.1 ISED Overview
Assuming $P$ is a black-box, we can collect symbol-output samples $(u,v)$ from $P$. Such collection of samples can be viewed as a summary logic program consisting of rules of the form if $r=u$ then $y=v$. For instance, in the task of adding two digits $r_{1}$ and $r_{2}$, one rule of the logic program would be $r_{1}=1\land r_{2}=2\rightarrow y=3$. Techniques from neurosymbolic literature via exact or approximate weighted model counting (WMC) [10] can then be used for computing the gradient across such a summary of $P$. However, having the complete summary of all combinations of symbols is not feasible for a black-box $P$. ISED samples symbols from the probability distribution predicted by the neural network $M_{\theta}$, evaluates $P$ on each sample, and takes the gradient across this partial summary of $P$. This is a good approximation of the complete summary since it is likely to contain symbols with high probability, which contribute the most in exact computation.
This approach differs from REINFORCE in how it differentiates through this summary of $P$. REINFORCE rewards sampled symbols that resulted in the correct output through optimizing the log probability of each symbol, weighted by reward values. This weighted-sum style estimation provides a weaker learning signal compared to WMC used by ISED, making learning harder for REINFORCE as the number of inputs to $P$ increases. See Appendix A for further details.
@@ 3.2 Preliminaries and Programming Interface
Using this, we may further define data-types such as $\textsc{Integer}_{j}^{k}=\textsc{Discrete}(\{j,\dots,k\})$, $\textsc{Digit}=\textsc{Integer}_{0}^{9}$, and $\textsc{Bool}=\textsc{Discrete}(\{\text{true},\text{false}\})$. These types give ISED the flexibility learn neural programs with diverse types of inputs and outputs, e.g., $\textsc{Permutation}_{n}$ input and output types for integer list sorting and $\textsc{List}_{9}(\textsc{List}_{9}(\textsc{Digit}))$ for sudoku solving.
We also define a black-box program $P$ as a function $(\tau_{1},\dots,\tau_{m})\rightarrow\tau_{o}$, where $\tau_{1},\dots,\tau_{m}$ are the input types and $\tau_{o}$ is the output type. For example, the structural input mapping for the hand-written formula task is $\textsc{List}_{7}(\textsc{Discrete}(\{0,\dots,9,+,-,\times,\div\}))$, and the structural output mapping is Float. The mappings suggest that the program takes a list of length up to 7 as input, where each element is a digit or an arithmetic operator, and returns a floating point number.
@@
In order to represent the ground truth output as a distribution to be used in the loss computation, there needs to be a mechanism for transforming SET($\tau$) mappings into DIST($\tau$) mappings. For this purpose, we define a vectorize function $\delta_{\tau}:(\text{SET}(\tau),2^{\tau})\rightarrow\text{DIST}(\tau)$ for the different output mappings $\tau$ in Table 2. When considering a datapoint $(x,y)$ during training, ISED samples many symbols and obtains a list of outputs $\hat{y}$. The vectorizer then takes the ground truth $y$ and the outputs $\hat{y}$ as input and returns the equivalent distribution interpretation of $y$. While $\hat{y}$ is not used by $\delta_{\tau}$ in most cases, we include it as an argument so that Float output mappings can be discretized, which is necessary for vectorization. For example, if the inputs to the vectorizer for the hand-written formula task are $y=2.0$ and $\hat{y}=[1.0,3.5,2.0,8.0]$, then it would return $[0,0,1,0]$.
@@ 3.3 Algorithm
We now formally present the ISED algorithm. For a given task, there is a black-box program $P$, taking $m$ inputs, that operates on structured data. Let $\tau_{1},...,\tau_{m}$ be the mappings for these inputs and $\tau_{o}$ the mapping for the program’s output. We write $P$ as a function from its input mappings to its output mapping: $P:(\tau_{1},...,\tau_{m})\rightarrow\tau_{o}$. For each unstructured input $i$ to the program, there is a neural model $M^{i}_{\theta_{i}}:x_{i}\rightarrow\text{DIST}(\tau_{i})$. $S$ is a sampling strategy (e.g., categorical sampling) that samples symbols using the outputs of a neural model, and $k$ is the number of samples to take for each training example. There is also a loss function $\mathcal{L}$ whose first and second arguments are the predicted and ground truth values respectively. We present the pseudocode of the algorithm in Algorithm 1 and describe its steps with the hand-written formula task:
@@
13:         $\hat{w}\leftarrow\text{normalize}([\omega(y_{l},\hat{y},\hat{r},\hat{p})\text{for }y_{l}\in\tau_{o}\text{ (or }y_{l}\in\hat{y})])$triangleright Estimate
@@
Estimate. In order to compute the prediction value to use in the loss function, ISED must consider each output $y_{l}$ in the output mapping and accumulate the aggregated probabilities for all sampled symbols that resulted in the output $y_{l}$. We specify $\otimes$ as the min function, and $\oplus$ as the max function in this example. Note that ISED requires that $\otimes$ and $\oplus$ represent either min and max or mult and add respectively. We refer to these two options as the min-max and add-mult semirings. We define an accumulate function $\omega$ that takes as input an element of the output mapping $y_{l}$, sampled outputs $\hat{y}$, sampled symbols $\hat{r}$, and predicted input distributions $\hat{p}$. The accumulator performs the $\oplus$ operation on aggregated probabilities for elements of $\hat{y}$ that are equal to $y_{l}$ and is defined as follows:
$\omega(y_{l},\hat{y},\hat{r},\hat{p})=\oplus_{j=1}^{k}\mathbf{1}_{\hat{y}_{j}=y_{l}}\sigma_{\tau_{o}}(\hat{r}_{j},\hat{p}_{j})$
Continuing our example, suppose, among the samples, there are two symbolic combinations ($[7,+,2]$ and $[3,*,3]$) that resulted in the output $9.0$. Let us say that these sets of symbols had probabilities $[0.3,0.8,0.8]$ and $[0.1,0.1,0.1]$, respectively. Then the result of the probability aggregation for $y_{l}=9.0$ would be $\omega(9.0,\hat{y},\hat{r},\hat{p})=\texttt{max}(\texttt{min}([0.3,0.8,0.8]),\texttt{min}([0.1,0.1,0.1]))=0.3$.
ISED then sets $\tilde{w}=[\omega(y_{l},\hat{y},\hat{r},\hat{p})\text{ for }y_{l}\in\tau_{o}]$ in the case where $\tau_{o}$ is not Float. When $\tau_{o}$ is Float, as for hand-written formula, it only considers $y_{l}\in\hat{y}$. Next, it performs $L_{2}$ normalization over each element in $\tilde{w}$ and sets $\hat{w}$ to this result. To initialize the ground truth vector, it sets $w=\delta(y,\hat{y})$. ISED then initializes $l=\mathcal{L}(\hat{w},w)$ and computes $\frac{\partial l}{\partial\theta_{i}}$ for each input $i$. These steps are shown on lines 13-16. In our running example, since $9.0$ is an incorrect output, the probability of the first symbol being equal to $7$ (instead of the correct answer $1$) will be penalized while the probabilities for predicting other symbols are unchanged.
@@ 4.1 Benchmark Tasks: NeuroGPT, NeuroPython, and Neurosymbolic
We also consider several tasks from the neurosymbolic literature, including hand-written formula (HWF) evaluation and Sudoku solving. While the solutions to many of these tasks are usually presented as a logic program in neurosymbolic learning frameworks, neural program solutions can take the form of Python programs. We call such models neuroPython programs.
@@
Visual Sudoku. The goal of this task is to solve an incomplete 9x9 Sudoku, where the problem board is given as MNIST digits. We follow the experimental setting of NASR [5], including their pre-trained MNIST digit recognition models and sudoku solvers. We use the SatNet dataset consisting of 9K training samples and 500 test samples [25].
@@ 4.2 Evaluation Setup and Baselines
We pick as baselines neurosymbolic methods DeepProbLog (DPL) [14] and Scallop [13], A-NeSI [24] which performs neural approximation of the gradients, and sampling-based gradient approximation methods REINFORCE [26], IndeCateR [21], and NASR [5]. IndeCateR achieves provably lower variance than REINFORCE by using a specialized sampling method (Appendix A), and NASR is a variant specialized for efficient finetuning by using a single sample and a custom reward function. We also use purely neural baselines and CLIP [19] for GPT leaf and scene. CLIP is a multimodal model that supports zero-shot image classification by simply providing names of the output categories.
@@ 4.3 RQ1: Performance and Accuracy
Table 3: Performance on selected benchmarks. "TO" means time-out, and "N/A" means the task could not be programmed in the framework. Methods are divided (from top to bottom) by neurosymbolic, black-box gradient estimation, and REINFORCE-based. [^3] In an earlier version of this paper, we reported lower numbers for REINFORCE and IndeCateR for RQ1 but have since updated these numbers to reflect their performance with the MNIST network from the IndeCateR implementation. Unlike ISED’s network, IndeCateR’s network lacks a softmax layer.
@@
Method sum2 sum3 sum4 HWF DT leaf GPT leaf scene sudoku
DPL $95.14$ $93.80$ TO TO $39.70$ N/A N/A TO
Scallop $91.18$ $91.86$ $80.10$ $96.65$ $81.13$ N/A N/A TO
A-NeSI $96.66$ $94.39$ $78.10$ $3.13$ $78.82$ $72.40$ $61.46$ $26.36$
REINFORCE $74.46$ $19.40$ $13.84$ $88.27$ $40.24$ $53.84$ $12.17$ $79.08$
IndeCateR $96.48$ $93.76$ $92.58$ $95.08$ $78.71$ $69.16$ $12.72$ $66.50$
NASR $6.08$ $5.48$ $4.86$ $1.85$ $16.41$ $17.32$ $2.02$ $82.78$
ISED (ours) $80.34$ $95.10$ $94.10$ $97.34$ $82.32$ $79.95$ $68.59$ $80.32$
To answer RQ1, we evaluate ISED’s accuracy against those of the baselines. ISED matches, and in many cases surpasses, the accuracy of neurosymbolic and gradient estimation baselines. We highlight the results for sumn from MNIST-R and other benchmarks in Table 3. Tables 7-10 in Appendix C contain results for the remaining MNIST-R tasks, including standard deviations for all tasks. ISED is the top performer on 8 out of the 16 total tasks.
On the GPT leaf and scene tasks, ISED outperforms the purely neural baseline by $3.82\%$ and $31.42\%$ respectively, and zero-shot CLIP by $59.80\%$ and $17.50\%$. For many tasks, A-NeSI is the non-neurosymbolic method that comes closest to ISED, sometimes outperforming our method. However, A-NeSI achieves significantly lower performance than ISED on tasks involving complex programs, namely HWF and sudoku. This is likely due to the difficulty of training a neural model to estimate the output of $P$ and its gradient when $P$ is complex. ISED also outperforms REINFORCE on all but 3 tasks due to the REINFORCE learning signal being weaker for tasks where $P$ involves multiple inputs. NASR outperforms ISED only on sudoku by $2.46\%$ due to NASR being well-suited for fine-tuning as it restricts its algorithm to use a single sample. IndeCateR achieves similar performance compared to ISED on most tasks but achieves significantly lower accuracy on the scene classification task, which has a large input space with maximum 10 objects each with 47 possible values in each scene, demonstrating that IndeCateR is less sample-efficient than ISED. We elaborate more on this point in RQ2.